An artifact was found and tested for its carbon-14 content. If 84% of the original carbon-14 was still present, what is its probable age (to the nearest 100 years)? Use that carbon-14 has half-life of 5,730 years.
_______ years?

Answer by stanbon(75887)

An artifact was found and tested for its carbon-14 content.
If 84% of the original carbon-14 was still present, what is its probable age (to the nearest 100 years)?
Use that carbon-14 has half-life of 5,730 years.
----
A(t) = Ao(1/2)^(t/5730)
----------
0.84*Ao = Ao(1/2)^(t/5730)
---
(1/2)^(t/5730) = 0.84
----
Take the log:
---
(t/5730) = log(0.84)/log(1/2) = 0.2515
---
t = 5730*0.2515 = 1441.32 years
========================================
Cheers,
Stan H.
==============